OpenStudy (anonymous):

B is twice as efficient as A and C is 50% more efficient than B. B and C together can complete a work in 10 days. How much time will be taken by A and B to complete the work, if they work on alternate days starting with A?

OpenStudy (phi):

This is an unusually complicated word problem

OpenStudy (anonymous):

Like really XD my brain goes so far and crashes

OpenStudy (phi):

Let's say A works at a rate of A jobs/day B is twice as efficient as A.... B= 2A (works twice as fast) C is 50% more efficient than B.... C= 1.5B= 3A (C works 3 times faster than A) B and C together can complete in 10 days: (B+C)*10 days = 1 job in terms of A's rate: (2A+3A)10=1 ... 50A= 1 A= 0.02 jobs/day (it takes A 50 days to complete one job)

OpenStudy (anonymous):

a:B:c=2:4:5, b+c=1/10 A:B:C 100:200:250 200+250=10so it is 45 then a is 100 b 200 ,300 300/450 is 3/2

OpenStudy (phi):

How much time will be taken by A and B to complete the work, if they work on alternate days starting with A? that means every 2 days you perform A+2A work: a rate of 3/2 A per day if they work an even # of days

OpenStudy (anonymous):

that i couldnt do it so only i am asking.for a pair of days it is2(a+b)=3/2 is 3/4

OpenStudy (phi):

working alternately, the work rate for A and B is 3/2 A per day, or 1.5*0.02= 0.03 jobs/day it will take 32 days to perform .96 of a job. (I need an *even* number) on the 33rd day A does 0.02... which gets us to 0.98 on the 34th day B could do 0.04... he only needs ½ of a day to do 0.02 so it takes A and B *33.5* days to complete the job. (it helps if I can count)

OpenStudy (anonymous):

the answer is correct or not how to check.is the method correct

OpenStudy (anonymous):

A is 25% more efficient than B. With the help of C, B can complete a work in border=0rd of the time which is required for A alone. How much percentage is C less efficient than A?

OpenStudy (phi):

Is this a different question? There are typos in this ** A is 25% more efficient than B. With the help of C, B can complete a work in border=0rd of the time ***

OpenStudy (anonymous):

A is 25% more efficient than B. With the help of C, B can complete a work in 2/3rd of the time which is required for A alone. How much percentage is C less efficient than A?

OpenStudy (anonymous):

A:B 125:100,5:4 , b+c=2/3a 3(b+c)=2A

OpenStudy (phi):

Here is how I think of these problems: rate * time = # of jobs rate is measured in jobs/day A is 25% more efficient than B means rate of A = 1.25 B (A works 25% faster) and B= 4/5 A = 0.8 A C, B can complete a work in 2/3rd of the time for A Let's say A takes T days: A*T= 1 job T= 1/A (B+C)* (⅔ T) = 1 job (4/5 A + C) * (⅔)*1/A = 1 job \[ \frac{4}{5}\cdot \frac{2}{3} \frac{A}{A} + \frac{2}{3}\frac{C}{A} = 1 \\ \frac{8}{15} + \frac{2}{3}\frac{C}{A} = 1 \\ \frac{2}{3}\frac{C}{A} = \frac{7}{15} \\ \frac{C}{A} =\frac{7}{15}\cdot \frac{3}{2}= \frac{21}{30} \] assuming I did not make a silly mistake.

OpenStudy (phi):

C's rate is 70% of A's C's rate is 30% less than A's

OpenStudy (anonymous):

wow ,fantastic @phi

OpenStudy (anonymous):

\[ \frac{4}{5}\cdot \frac{2}{3} \frac{A}{A} + \frac{2}{3}\frac{C}{A} = 1 \\ \frac{8}{15} + \frac{2}{3}\frac{C}{A} = 1 \\ \frac{2}{3}\frac{C}{A} = \frac{7}{15} \\ \frac{C}{A} =\frac{7}{15}\cdot \frac{3}{2}= \frac{21}{30} \] what is this?. i cant understand

OpenStudy (phi):

They say C, B can complete a work in 2/3rd of the time for A I let T = the time it takes A to do 1 job... so B and C together will take ⅔ T days The equation is rate* time = jobs done (similar to rate*time = distance ) in other words C, B can complete a work in 2/3rd of the time for A becomes (B+C)*⅔ T = 1 job

OpenStudy (anonymous):

i understand that but only fraction/

OpenStudy (phi):

but we want things in terms of A (not B) we know A= 1.25B so B= 4/5 A we know A*T= 1 job so T= 1/A use those in the equation and solve for C/A (ratio of C's rate to A's rate)

OpenStudy (phi):

Does this make sense (as an example) X can do a job in 10 days. means X does 1/10 of a job every day. if X works for 10 days he does 1/10 job/day * 10 days= 1 job if X works for 5 days he does 1/10 job/day * 5 days= ½ job if X works for 20 days he does 1/10 job/day * 20 days= 2 jobs

OpenStudy (phi):

if Y also does a job in 10 days, Y does 1/10 of a job per day X and Y together do 1/10 + 1/10 = 2/10= 1/5 job in 1 day we can write: (1/10 + 1/10)*T = # of jobs done by X and Y together in T days

OpenStudy (anonymous):

k i understood @phi .i asked only about/{fraction like that why it is error

OpenStudy (phi):

try refreshing your browser

OpenStudy (anonymous):

k

OpenStudy (phi):

It should look like this